A gas at STP occupies 28cm^3 of space. If the pressure changes to 3.8 ATM and the temp increases at 203 degrees celcius, did the new volume. (Hint: 1cm^3 = 1 mL)
1
Expert's answer
2013-03-13T06:54:15-0400
STP: pressure - 105 Pa = 0.98692 ATM; temperature - 0 C (273 К).
This task can be solved by such equation:
pV = nRT
V = nRT/p
All the data except T and p is constant. So if the ratio between T and p remains constant, the volume will not change.
1) 273/0.98692 = 276.6
2) (203 + 273)/3.8 = 125.26
So the volume will become less in 2.2 times: 28 ml/2 = 14 ml